Commander Report – December 7th

Sol: 10
Summary Title: Snow Day and EVA #14
Author’s name: Dr. Kristine Ferrone, Commander
Mission Status: Nominal
Commander’s Report:

Crew 269 awoke to a winter wonderland landscape this morning and snowy conditions. The crew enjoyed a quiche bake for breakfast and conducted a morning meeting to discuss how the weather would affect the day’s EVA plans. After a consultation with Tier 1 mission support, the crew decided to shorten the morning EVA #14 and continue to send 2 crew to perform repair operations on the mirror coating experiment payload. HSO and Engineer traversed on foot through the muddy terrain to repair the mirror payload and return swiftly to the Hab, while Green Hab Officer lead IVA comms. During the EVA, CDR and XO discussed end of mission reporting requirements and began to create content for those reports while Technology Officer continued to work data processing on the Aerospace computers. In the afternoon, HSO collected health and safety stats, crew completed exercise protocols, and Green Hab Officer tended to the plants in the Green Hab. Power was again limited due to weather conditions and the crew followed Tier 1 support instructions to conserve power. For dinner, the crew enjoyed black bean burgers on homemade sourdough buns with microgreens from the Green Hab.

Copyright © The Mars Society. All rights reserved. | Main Site